OpenStudy (anonymous):

What are the solutions for x in the proportion 5x-6/x=2x/x ? x = −12 and x = −8 x = 12 and x = 8 x = −2 and x = −3 x = 2 and x = 3

OpenStudy (anonymous):

How do you solve a proportion?

OpenStudy (anonymous):

cross multipy. which would give me 5x-6(x)=2x(x)

OpenStudy (anonymous):

Careful. (5x) (x) = (2x)(x) =

OpenStudy (anonymous):

5x^2 & 2X^2?

OpenStudy (anonymous):

Right. Now gather like terms and solve. If you are left with x^2 terms gather everything on one side and factor. What do YOU get?

OpenStudy (anonymous):

This is always the part i don't understand. I didn't get it in the lesson. can you explain to me how to actually find the answers after all of that?

OpenStudy (anonymous):

5x^2 -6x = 2x^2 Gather x^2 on same side. 3x^2 -6x = 0 Agree?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Get a GCF for the two terms and factor it outside the parentheses. Whadda ya git?

OpenStudy (anonymous):

45

OpenStudy (anonymous):

GCF=3 but i don't see how i would factor that... 3[(x^2-2x)]

OpenStudy (anonymous):

Now we can also factor out x. 3x ( x -2). What x makes 3x=0? What x makes (x-2) = 0? These are possible answers. You have to check to be sure they work. What happened now?

OpenStudy (anonymous):

subtract the 3 and add the 2?

OpenStudy (anonymous):

2 is good, but 3 times what =0

OpenStudy (anonymous):

0

OpenStudy (anonymous):

So we have two possibilities 2 and 0. Do both work in the original proportion?

OpenStudy (anonymous):

it doesn't look lik it to me. i might have done my work wrong. plus 0 isnt an answer.

OpenStudy (anonymous):

0 would cause division by 0, so it is no good. 2 definitely works, but 3 will not work even though it is part of the same answer as 2. Peculiar???

OpenStudy (anonymous):

Very

OpenStudy (anonymous):

My fancy calculator confirms OUR answer.